OpenStudy (anonymous):

Consider the following pair of equations: y = x + 4 y = −2x − 2 Can someone please help me with this? Will fan and give medal. Explain how you will solve the pair of equations by substitution.

OpenStudy (triciaal):

y = y example 2 + 8 = 10 and 6 + 4 = 10 therefore 2 + 8 = 6 + 4 set the expressions equal and solve for x put this value of x in either expression for y and find the value of y good practice to check your solution in the original because it is very easy and very common to make arithmetic errors

OpenStudy (triciaal):